The perimeter of a triangle is 6x^2+11 inches. If two sides have lengths of 3x+2 and x^2+6, find the polynomial that represents the third side of the triangle. Explain your reasoning.
Answer by rwm(914)

call the perimeter P=6x^2+11
one side is a=3x+2
other side is b=x^2+6
missing side is c=P-a-b
just do the subtraction
since P=a+b+c
we want c
c=P-a-b
c=6x^2+11-(3x+2)-(x^2+6)
c=6x^2+11-3x-2-x^2-6
c=5x^2-3x+3